If the nominal exchange rate is 12 Mexican pesos per $1Question 9 (1 point) If the nominal exchange rate is 12 Mexican pesos per $1, a margarita that costs 36 pesos in Mexico costs
Answer #1

12 mexian pesos = 1 Dollar

1 mexian pesos = 1/12 Dollar

36 mexian pesos = 36 * 1/12 Dollar = 3 dollar

Last option is correct answer